mr. salazar bought 3 shirts and 2 pants for php.2150. a month later , at the same price, he bought 2 shirts and 4 pants for php.2900. what is the cost of a shirt?
Answer by mananth(16946)

shirt price x
pant price y
..
3x+2y=2150
2x+4y=2900
multiply (1)by -2
Multiply (2) by 3
-6x-4y=-4300
6x+12y=8700

Add the two equations
8y=4400
/8
y=550 pant
Plug value of y in
3x+2y=2150
3x+1100 =2150
3x=1050
/3
x=350 shirt